### v1

# Instabilities of Gauged Q-Balls

Abstract
We present the first analytical calculation that shows that perturbations with angular dependence can lead to an instability in gauged Q-balls. We find an explicit condition on the parameters for the Q-ball to become unstable. We compare our predictions to the numerical calculation in Kinach et al., and show agreement, including a correct prediction of the instability/stability of the two parameter points analyzed in that paper.
preprint: UCI-TR-2024-08
I Introduction
In certain scalar field theories, one can construct solitonic solutions called Q-balls Coleman:1985ki . For these solutions to exist, the scalar field must have a $U(1)$ symmetry, and in addition, the potential must satisfy certain conditions (we review this in Sec. II). If the symmetry is ungauged, the resulting Q-balls are called global Q-balls, while if the symmetry is gauged, they are called gauged Q-balls Lee:1988ag ; Gulamov:2013cra ; Gulamov:2015fya ; Nugaev:2019vru .
The stability of these solitons has been discussed in many papers, and the picture is still unclear. For global Q-balls, it has been shown Lee:1991ax ; Friedberg:1976me ; Paccetti:2001uh that the Q-balls are stable as long as ${d\omega\over dQ}<0$ (these quantities are defined below). For gauged Q-balls, the analogous statement is still believed to be true, but this has not been proven Panin:2016ooo .
Recently, a numerical analysis of the stability of gauged Q-balls was carried out Kinach:2022jdx . It was found that gauged Q-balls were unstable for certain parameters, and in particular were unstable to modes which had an angular dependence; indeed, for one particular parameter choice that was studied (denoted P2 in that paper), the instability had $L=4$. On the other hand, for other choices of parameters, the corresponding Q-balls were stable. In general, Q-balls with ${d\omega\over dQ}<0$ were stable.
Our goal here is to provide an analytic approach to studying these instabilities. This will be based on the approach to Q-balls developed in Heeck:2020bau ; Heeck:2021zvk . These papers showed that in the large radius limit, spherical Q-balls can be analytically described by considering corrections around the thin wall limit (this will be reviewed below in Sec. III). By considering a similar analysis for slightly nonspherical gauged Q-balls, we find a possible mode of instability for these Q-balls. This analysis is in section IV.
We then show in section V that the instability that we find has many characteristics in common with the numerical analysis of Kinach:2022jdx . In particular, we indeed find that for the parameter choice P2 in Kinach:2022jdx , the $L=4$ mode is unstable. Similarly, we find that for another parameter choice P1, the instability is absent, in agreement with the numerical analysis. We also find that Q-balls with ${d\omega\over dQ}<0$ are stable.
We close with comments and open questions.
II Q-Balls: Overview and Notation
We review here the basic properties of Q-balls; we will mostly follow the notation of Heeck:2020bau ; Heeck:2021zvk .
The Lagrangian density for a charged scalar field is
$\mathcal{L}=\left|D_{\mu}\phi\right|^{2}-U(|\phi|)-\frac{1}{4}F_{\mu\nu}F^{\mu\nu},$ (1)
where $D_{\mu}=\partial_{\mu}-\text{i}eA_{\mu}$ is the gauge covariant derivative and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the field-strength tensor. Q-balls exist when the function $U(|\phi|)/|\phi|^{2}$ has a minimum at $0<\phi_{0}<\infty$ such that
$0\leq\sqrt{\frac{2U(\phi_{0})}{\phi_{0}^{2}}}\equiv\omega_{0}<m_{\phi}\,,$ (2)
where $m_{\phi}$ is the mass of the complex scalar. We make the static charge ansatz Lee:1988ag
$\phi(t,\vec{x})=\frac{\phi_{0}}{\sqrt{2}}f(r)e^{\text{i}\,\omega t}\,,\qquad A_{0}(t,\vec{x})=A_{0}(r)\,,\qquad A_{i}(t,\vec{x})=0\,,$ (3)
The scalar frequency $\omega$ is restricted to the region $\omega_{0}<\omega\leq m_{\phi}$.
It is convenient to define the dimensionless quantities
$\rho\equiv r\sqrt{m_{\phi}^{2}-\omega_{0}^{2}},\quad A(\rho)\equiv\frac{A_{0}(\rho)}{\phi_{0}}\,,\quad\alpha\equiv e\Phi_{0}\,,\quad\{\Omega_{0},\Omega,\Phi_{0}\}\equiv\frac{\{\omega_{0},\omega,\phi_{0}\}}{\sqrt{m_{\phi}^{2}-\omega_{0}^{2}}}$ (4)
The equations of motion are
$\nabla^{2}f$ $=\frac{1}{\Phi_{0}^{2}(m_{\phi}^{2}-\omega_{0}^{2})^{2}}\frac{\text{d}U}{\text{d}f}-\left(\Omega-\alpha A\right)^{2}f$ (5)
$\nabla^{2}A$ $=\alpha f^{2}(A\alpha-\Omega)$ (6)
The charge is
$Q=\Phi_{0}^{2}\int d^{3}xf^{2}\left(\Omega-\alpha A\right)$ (7)
while the energy is
$E/\sqrt{m_{\phi}^{2}-\omega_{0}^{2}}=\Phi_{0}^{2}\int d^{3}x\left\{\frac{1}{2}f^{\prime 2}+\frac{1}{2}A^{\prime 2}+\frac{1}{2}f^{2}\left(\Omega-\alpha A\right)^{2}+\frac{U(f)}{\Phi_{0}^{2}(m_{\phi}^{2}-\omega_{0}^{2})^{2}}\right\}$ (8)
For concreteness we restrict most of our discussion to the sextic scalar potential studied in Ref. Heeck:2020bau . This can be parametrized as
$U(f)=\phi_{0}^{2}\left(\frac{m_{\phi}^{2}-\omega_{0}^{2}}{2}\,f^{2}(1-f^{2})^{2}+\frac{\omega_{0}^{2}}{2}\,f^{2}\right).$ (9)
III Spherical gauged Q-balls
This section follows the analysis of Heeck:2021zvk .
Large spherical Q-balls are described by a scalar profile that is approximately a step function;
$f^{(0)}\to\left\{\begin{array}[]{c}f_{+}\qquad\rho<R\\ 0\qquad\rho>R\end{array}\right.$ (12)
More precisely, we can divide the Q-ball into three regions.
In the interior $\rho\ll R$, the scalar is extremely close to a solution of the RHS of (5) i.e. $f^{(0)}\sim f_{+}$ where
$\left(\frac{1}{\Phi_{0}^{2}(m_{\phi}^{2}-\omega_{0}^{2})^{2}}\frac{\text{d}U}{\text{d}f}-\left(\Omega-\alpha A\right)^{2}f\right)_{{f=f_{+}}}=0.$ (13)
For the specific potential (9), we have $f_{+}\sim 1$.
In this region, the gauge field is massive and satisfies
$\nabla^{2}A=\alpha f_{+}^{2}(A\alpha-\Omega)$ (14)
We will take $f_{+}$ to be approximately constant. In that case, the solution for the gauge field is
$\alpha A^{(0)}=\Omega+Q_{in}{\sinh(\alpha f_{+}\rho)\over\rho}$ (15)
In the exterior $\rho\gg R$, the scalar is extremely close to zero. In this region, the solution for the gauge field is
$\alpha A^{(0)}={Q_{out}\over\rho}$ (16)
Continuity of the field and its derivative at $\rho=R$ then implies
$Q_{in}=-{\Omega\over{\alpha f_{+}\cosh(\alpha f_{+}R)}}\qquad\qquad Q_{out}=R\Omega\left(1-{1\over{\alpha f_{+}R}}{\tanh(\alpha f_{+}R)}\right)$ (17)
The transition region $\rho\sim R$ is the most interesting. In this region, the scalar varies quickly from $f_{+}$ to zero over a range of order 1. The gauge field on the other hand varies much more slowly (for small $\alpha$) and can be taken to be the matching value of the solution (15, 16, 17). We therefore take the gauge field to be approximately constant with the value
$\alpha A_{R}^{(0)}=\Omega\left(1-{1\over\alpha f_{+}R}{\tanh(\alpha f_{+}R)}\right)$ (18)
The equation for $f$ is now
$(f^{(0)})^{\prime\prime}+\frac{2}{R}(f^{(0)})^{\prime}=f^{(0)}\left(1-4(f^{(0)})^{2}+3(f^{(0)})^{4}+{\Omega_{0}^{2}}-(\Omega_{R}^{(0)})^{2}\right)$ (19)
where we have replaced the explicit factor of $\rho$ by $R$, and
$\Omega_{R}^{(0)}=\left(\Omega-\alpha A^{(0)}_{R}\right)={\Omega~{}{\tanh(\alpha f_{+}R)}\over\alpha f_{+}R}$ (20)
Since $\Omega_{R}^{(0)}$ is being treated as effectively constant, this maps the problem to a global Q-ball equation where $\Omega_{R}^{(0)}$ plays the role of the frequency. Since the frequency and the radius are related for a global Q-ball, this equation can only be solved if $\Omega_{R}^{(0)}=\Omega_{e=0}$, where $\Omega_{e=0}$ is the value of $\Omega$ which would produce a global Q-ball of radius $R$.
Following Heeck:2021zvk , the solution of the equation (19) is
$f^{(0)}(\rho)=\frac{f_{+}}{\sqrt{1+2e^{2(\rho-R)}}}\,.$ (21)
with
$((\Omega_{R}^{(0)})^{2}-\Omega_{0}^{2})R=1+{\cal O}({1\over R})$ (22)
This can be shown to be an accurate description of the Q-ball profile Heeck:2021zvk .
IV Perturbed Q-balls
IV.1 Ansatz for the perturbation
We now consider a perturbation of the Q-ball where the surface becomes nonspherical; we take the surface to be at
$\rho=R(\theta)\equiv R(1+\epsilon P_{L}(\cos(\theta))).$ (23)
We will below, for conciseness, denote the Legendre polynomials $P_{L}(\cos(\theta))$ as $P_{L}$.
We modify the transition profile to reflect the perturbation; we can write the ansatz for the new profile as
$f(\rho)=\frac{f_{+}}{\sqrt{1+2e^{2(\rho-R(\theta))}}}\,.$ (24)
Typically, such a perturbation will not lead to a solution of the equations of motion, and the perturbation will either grow or decay. In either case, $\epsilon$ will have a time dependence. Even if we have a perturbation which is initially static (i.e. the initial time derivative $\dot{\epsilon}=0$), the dynamics will drive the value of $\epsilon$ away from its initial value. That is, the dynamics induces a nonzero value of $\ddot{\epsilon}$. Instead of a constant $\epsilon$, therefore, we can set (for small times) $\epsilon(t)=\epsilon(1+{1\over 2}\gamma t^{2})$. If $\gamma<0$, the perturbation is presumably stable, while if $\gamma>0$, the perturbation grows, and therefore is unstable. We will assume however that $\gamma$ is small, so that we can approximately treat $\epsilon$ as constant.
We now look for an equation that determines $\gamma$.
IV.2 The gauge field
The first step is to calculate the profile of the gauge field, and in particular the value of the gauge field in the transition region.
Following the previous section, we use the thin wall approximation to find the gauge field. In the interior we will again have $f=f_{+}$ while in the exterior $f=0$. The equations of motion for the gauge field in the interior and exterior are therefore unchanged.
At the linearized level, the gauge field will obtain harmonics proportional to $P_{L}$. The solution will be therefore of the form
$\alpha A=\left\{\begin{array}[]{c}\Omega+Q_{in}{\sinh(\alpha f\rho)\over\rho}+\epsilon P_{in}F_{L}(\alpha f\rho)P_{L}\qquad\rho\ll R(\theta)\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}Q_{out}{1\over\rho}+\epsilon P_{out}{1\over\rho^{L+1}}P_{L}\qquad~{}~{}~{}~{}~{}~{}~{}~{}\rho\gg R(\theta)\end{array}\right.$ (27)
where $F_{L}(\alpha f\rho)=j_{L}(i\alpha f\rho)$, and $j_{L}$ is a spherical Bessel function. We list the first few $F_{L}$ in an appendix.
We again require the gauge field and its derivative to be continuous at the boundary $\rho=R(\theta)$. This yields
$P_{in}={\Omega\tanh(\alpha fR)\over\alpha fF^{\prime}_{L}(\alpha f\rho)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}\qquad\qquad P_{out}={\Omega R^{L+1}\tanh(\alpha fR)F_{L}(\alpha fR)\over F^{\prime}_{L}(\alpha fR)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}$ (28)
This then implies that at $\rho=R(\theta)$, we have
$\alpha A_{R}=\alpha A^{(0)}_{R}+\epsilon P_{L}a_{R}$ (29)
where
$a_{R}={\Omega F_{L}(\alpha fR)\tanh(\alpha fR)\over F^{\prime}_{L}(\alpha f\rho)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}-\Omega\left(1-{\tanh(\alpha fR)\over\alpha fR}\right)$ (30)
IV.3 The scalar equation
Once again we note that the gauge field is slowly varying as a function of $\rho$, and can be taken to have a value $A_{R}(\theta)$ over the transition region. We then have
$-\ddot{f}+f^{\prime\prime}+\frac{2}{\rho}f^{\prime}+{1\over\rho^{2}\sin\theta}\partial_{\theta}(\sin\theta\partial_{\theta}f)=f\left(1-4f^{2}+3f^{4}+{\Omega_{0}^{2}}-\Omega_{R}^{2}\right)$ (31)
where $\Omega_{R}=\Omega-\alpha A_{R}(\theta)$.
In the ansatz (24), the time dependence and the angular dependence are linked to the radial dependence. Indeed, we can write
$-\ddot{f_{T}}={f_{T}}^{\prime}R\gamma{\epsilon}P_{L}$ (32)
and similarly
${1\over\rho^{2}\sin\theta}\partial_{\theta}(\sin\theta\partial_{\theta}f_{T})=f_{T}^{\prime}(\epsilon){1\over R}L(L+1)P_{L}$ (33)
Note that this is small, and so to leading order we can treat the scalar field as only depending on $\rho$.
Substituting these into equation (31), we find that in the transition region, the equation is
$f^{\prime\prime}+\frac{2}{R_{eff}}f^{\prime}=f\left(1-4f^{2}+3f^{4}+{\Omega_{0}^{2}}-\Omega_{R}^{2}\right)$ (34)
with
$\frac{2}{R_{eff}}=\frac{2}{R(\theta)}+R\gamma\epsilon P_{L}+\epsilon{1\over R}L(L+1)P_{L}$ (35)
The equation (34) is identical to the equation (19)in the previous section with the replacement $\Omega^{(0)}_{R}\to\Omega_{R},R^{(0)}\to R_{eff}$ The relation (22) is therefore modified at large radius to
$\Omega_{R}^{2}-{\Omega_{0}^{2}}={1\over R_{eff}}+{\cal O}({1\over R^{2}_{eff}})$ (36)
Eliminating $R_{eff}$, we find
$2(\Omega_{R}^{2}-{\Omega_{0}^{2}})=2((\Omega^{(0)}_{R})^{2}-{\Omega_{0}^{2}})+R\gamma\epsilon P_{L}+\epsilon{1\over R}(L(L+1)-2)P_{L}$ (37)
This implies
$-4\Omega^{(0)}_{R}a_{R}-{(L+2)(L-1)\over R}=R\gamma$ (38)
where $a_{R}$ is defined in (29).
IV.4 The prediction for the instability
Combining (30) and (38), we have
$R^{0}\gamma=-4\Omega\Omega^{(0)}_{R}\left({F_{L}(\alpha fR)\tanh(\alpha fR)\over F^{\prime}_{L}(\alpha f\rho)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}-\left(1-{\tanh(\alpha fR)\over\alpha fR}\right)\right)-{(L+2)(L-1)\over R}$ (39)
where
$\Omega=\Omega_{R}^{(0)}{\alpha f_{+}R\over{\tanh(\alpha f_{+}R)}}\qquad\Omega^{(0)}_{R}=\left(\sqrt{\Omega_{0}^{2}+{1\over R^{2}}}\right)$ (40)
This is our final result for $\gamma$ in terms of the other parameters. As we have said above, $\gamma>0$ indicates an instability.
V Comparison with Numerics
We now compare to the numerical results of  Kinach:2022jdx .
V.1 The point P2
The authors of Kinach:2022jdx consider the sextic potential with
$V=|\phi|^{2}-(1/2)|\phi|^{4}+{0.2\over 3}|\phi|^{6}$ (41)
They analyzed in detail a specific parameter choice point (which they denote as P2), with $\omega=0.9958,e=0.02,r=65$ in their conventions. After rescaling to match our conventions, we find that this parameter point corresponds to
$\Omega_{0}^{2}=1/15\qquad\Omega=1.028\qquad R=62.9\qquad\alpha=0.057\qquad\alpha R=3.56$ (42)
From (40), we also have
$\Omega^{(0)}_{R}=\left(\sqrt{1/15+{1\over 62.9^{2}}}\right)=0.2587$ (43)
We now calculate the value of $R\gamma$ for each $L$. The results are shown in Table 1.
L 1 2 3 4 5 6
$R\gamma$ 0 0.123 0.148 0.102 -0.0001 -0.15
Table 1: $R\gamma$ as a function of $L$ for the point P2
We first note that $L=1$ appears to be a mode which is marginal i.e. neither stable nor unstable. This mode in fact corresponds to a translation of the Q-ball, and will be henceforth ignored.
We see from the above that $L=2,3,4$ all have $\gamma>0$, and hence are all predicted to be unstable modes. $L=5$ is apparently stable, but the value is very close to zero, and this result may be changed in a more accurate analysis. Other modes are stable.
This is to be compared with the results of Kinach:2022jdx , which finds that the Q-ball with this parameter choice is unstable, with an instability mode with $L=4$. They also find the instability is near the surface of the Q-ball. This is all consistent with our results. The only discrepancy is that we predict an $L=3$ mode which is in some sense more unstable; however, the absence of this mode, while somewhat puzzling, may plausibly be attributed to some feature of the numerical analysis.
V.2 Location of the Instability Threshold
We now look at the predicted instability threshold for $\alpha=0.057$, i.e. the first value of $R$ for which some $\gamma$ becomes positive. To do this, we calculate $R\gamma$ as a function of $R$. For $\Omega^{(0)}_{R}$, we use
$\Omega^{(0)}_{R}=\left(\sqrt{1/15+{1\over R^{2}}}\right)$ (44)
For $\Omega$, we scale this from the point P2 as
${\Omega\over 1.028}={\Omega_{R}^{(0)}\over 0.2587}{\alpha f_{+}R\over{\tanh(\alpha f_{+}R)}}{\tanh(3.26)\over 3.26}$ (45)
We then find that the smallest radius of instability occurs for $L=2$, and for a predicted value of $\alpha R=1.99$. We therefore predict that all radii above this value are unstable.
We compare this to the radius where ${d\Omega\over dR}$ becomes positive. This is at $\alpha R=0.51$, and it can then be checked straightforwardly, that for the radii which are in the unstable regime, we indeed have ${d\Omega\over dR}>0$.
However, it is to be noted that here we predict a significant difference between the instability threshold and the location where ${d\Omega\over dR}>0$. This is not in agreement with the results of Kinach:2022jdx , where these two appear to be at very similar values. We discuss this further in the next section.
V.3 Larger couplings: the point P1
The second parameter point discussed in Kinach:2022jdx is denoted P1, and has a much larger value of $e=0.17$, corresponding to a value of $\alpha=0.481$.
For such large values of $\alpha$, the analysis of the previous section cannot be trusted. Nevertheless, we can see what our formula predicts. Following the same procedure as the previous subsection, we find that the instability threshold is predicted to occur for $L=2$ at a value $\alpha R=6.7$. This corresponds to a predicted threshold radius of about 14. The Q-ball radius is not given in Kinach:2022jdx , but general considerations Loginov:2020xoj show that $R$ should scale as $\alpha^{-1}$ and should therefore be about 8. This indicates that the instability has not yet been reached, and P1 is stable. This is in agreement with the numerical results.
V.4 Smaller couplings
On the other hand, we can ask what happens if the coupling is taken to be smaller. We find that the value of $\alpha R$ at the threshold decreases. For example if we take $\alpha=10^{-4}$, we find a threshold radius (i.e. the radius where $\gamma=0$) that satisfies $\alpha R=0.21$. Numerical analysis of these very small $\alpha$ will shed light on whether our formulae are accurate for all Q-balls.
VI Discussion and Conclusion
We have constructed an explicit perturbation of a gauged Q-ball, and shown that in some parameter ranges, the equations of motion directly show that these perturbations can be unstable. These perturbations have angular dependence, and are localized near the surface of the Q-ball. We have found an explicit prediction for the condition on the radius (and other parameters) for the modes to become unstable.
This is, to our knowledge, the first calculation that shows analytically that perturbations with angular dependence can lead to an instability in gauged Q-balls.
We have compared our predictions to the numerical calculation in Kinach:2022jdx , and shown that we correctly predict the instability/stability of the two parameter points analyzed in that paper. Furthermore, it was found in the numerical analysis that one of the points had an unstable mode with $L=4$, and our formula is able to reproduce this by showing that the $L=4$ mode is indeed unstable. Another parameter point was found to be stable, and this is reproduced by our formula. We also find that increasing the gauge coupling makes the Q-balls more stable, a surprising result which was also found in the numerics.
We have also used our formula to characterize the boundary between stable and unstable parameters. This boundary does not agree perfectly with the results of Kinach:2022jdx . It appears that any Q-ball found to be unstable by our formula is also found to be unstable in the numerical calculations; however, there appear to be Q-balls which are found numerically to be unstable which are not captured by our analysis. This suggests strongly that there are other perturbations that we have missed that will extend the region of instability. On a related note, the region of instability in the numerical results appears to be linked to the condition ${d\omega\over dQ}>0$, at least for small gauge couplings. We do not see such a link, which is again suggestive that a different form of perturbation may be relevant for Q-balls. This analysis is left for future work.
VII Acknowledgments
This work was supported in part by NSF Grant No. PHY-2210283.
VIII Appendix A: Radial solutions
For completeness, we list the first few $F_{L}$ solutions used in the paper
$F_{0}$ $={\sinh x\over x}$ (46)
$F_{1}$ $=\sinh x\left({1\over x^{2}}\right)-\cosh x\left({1\over x}\right)$ (47)
$F_{2}$ $=-\cosh x\left({3\over x^{2}}\right)+\sinh x\left({1\over x}+{3\over x^{3}}\right)$ (48)
$F_{3}$ $=-\cosh x\left({1\over x}+{15\over x^{3}}\right)+\sinh x\left({6\over x^{2}}+{15\over x^{4}}\right)$ (49)
$F_{4}$ $=-\cosh x\left({10\over x^{2}}+{105\over x^{4}}\right)+\sinh x\left({1\over x}+{45\over x^{3}}+{105\over x^{5}}\right)$ (50)
$F_{5}$ $=-\cosh x\left({1\over x}+{105\over x^{3}}+{945\over x^{5}}\right)+\sinh x\left({15\over x^{2}}+{420\over x^{4}}+{945\over x^{6}}\right)$ (51)
$F_{6}$ $=-\cosh x\left({21\over x^{2}}+{1260\over x^{4}}+{10395\over x^{6}}\right)+\sinh x\left({1\over x}+{210\over x^{3}}+{4725\over x^{5}}+{10395\over x^{7}}\right)$ (52)