### v3

# Instabilities of Gauged Q-Balls

Abstract
We present the first analytical calculation that shows that perturbations with angular dependence can lead to an instability in gauged Q-balls. We find an explicit condition on the parameters for the Q-ball to become unstable. We compare our predictions to the numerical calculation in Kinach et al., and show agreement, including a correct prediction of the instability/stability of the two parameter points analyzed in that paper.
preprint: UCI-TR-2024-08
I Introduction
In certain scalar field theories, one can construct solitonic solutions called Q-balls Coleman:1985ki . For these solutions to exist, the scalar field must have a $U(1)$ symmetry, and in addition, the potential must satisfy certain conditions (we review this in Sec. II). If the symmetry is ungauged, the resulting Q-balls are called global Q-balls, while if the symmetry is gauged, they are called gauged Q-balls Lee:1988ag ; Gulamov:2013cra ; Gulamov:2015fya ; Nugaev:2019vru .
The stability of these solitons has been discussed in many papers, and the picture is still unclear. For global Q-balls, it has been shown Lee:1991ax ; Friedberg:1976me ; Paccetti:2001uh that the Q-balls are stable as long as ${d\omega\over dQ}<0$ (these quantities are defined below). For gauged Q-balls, the analogous statement is known to not be true Panin:2016ooo .
Recently, a numerical analysis of the stability of gauged Q-balls was carried out Kinach:2022jdx . It was found that gauged Q-balls were unstable for certain parameters, and in particular were unstable to modes which had an angular dependence. On the other hand, for other choices of parameters, the corresponding Q-balls were stable. In general, Q-balls with ${d\omega\over dQ}<0$ were stable.
Our goal here is to provide an analytic approach to studying these instabilities. This will be based on the approach to Q-balls developed in Heeck:2020bau ; Heeck:2021zvk . These papers showed that in the large radius limit, spherical Q-balls can be analytically described by considering corrections around the thin wall limit (this will be reviewed below in Sec. III). By considering a similar analysis for slightly nonspherical gauged Q-balls, we find a possible mode of instability for these Q-balls. This analysis is in section IV.
We then show in section V that the instability that we predict agrees in many details in common with the numerical analysis of Kinach:2022jdx . In particular, we indeed find that for one parameter choice in Kinach:2022jdx , the $L=4$ mode is unstable, in agreement with the numerical results. Similarly, we find that for another parameter choice, the instability is absent, in agreement with the numerical analysis.
We close with comments and open questions.
II Q-Balls: Overview and Notation
We review here the basic properties of Q-balls; we will follow the notation of Heeck:2020bau ; Heeck:2021zvk .
The Lagrangian density for a charged scalar field is
$\mathcal{L}=\left|D_{\mu}\phi\right|^{2}-U(|\phi|)-\frac{1}{4}F_{\mu\nu}F^{\mu\nu},$ (1)
where $D_{\mu}=\partial_{\mu}-\text{i}eA_{\mu}$ is the gauge covariant derivative and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the field-strength tensor. Q-balls exist when the function $U(|\phi|)/|\phi|^{2}$ has a minimum at $0<\phi_{0}<\infty$ such that
$0\leq\sqrt{\frac{2U(\phi_{0})}{\phi_{0}^{2}}}\equiv\omega_{0}<m_{\phi}\,,$ (2)
where $m_{\phi}$ is the mass of the complex scalar. We make the static charge ansatz Lee:1988ag
$\phi(t,\vec{x})=\frac{\phi_{0}}{\sqrt{2}}f(r)e^{\text{i}\,\omega t}\,,\qquad A_{0}(t,\vec{x})=A_{0}(r)\,,\qquad A_{i}(t,\vec{x})=0\,,$ (3)
The scalar frequency $\omega$ is restricted to the region $\omega_{0}<\omega\leq m_{\phi}$.
It is convenient to define the dimensionless quantities
$\rho\equiv r\sqrt{m_{\phi}^{2}-\omega_{0}^{2}},\quad A(\rho)\equiv\frac{A_{0}(\rho)}{\phi_{0}}\,,\quad\alpha\equiv e\Phi_{0}\,,\quad\{\Omega_{0},\Omega,\Phi_{0}\}\equiv\frac{\{\omega_{0},\omega,\phi_{0}\}}{\sqrt{m_{\phi}^{2}-\omega_{0}^{2}}}$ (4)
The equations of motion are
$\nabla^{2}f$ $=\frac{1}{\Phi_{0}^{2}(m_{\phi}^{2}-\omega_{0}^{2})^{2}}\frac{\text{d}U}{\text{d}f}-\left(\Omega-\alpha A\right)^{2}f$ (5)
$\nabla^{2}A$ $=\alpha f^{2}(A\alpha-\Omega)$ (6)
The charge is
$Q=\Phi_{0}^{2}\int d^{3}xf^{2}\left(\Omega-\alpha A\right)$ (7)
while the energy is
$E/\sqrt{m_{\phi}^{2}-\omega_{0}^{2}}=\Phi_{0}^{2}\int d^{3}x\left\{\frac{1}{2}f^{\prime 2}+\frac{1}{2}A^{\prime 2}+\frac{1}{2}f^{2}\left(\Omega-\alpha A\right)^{2}+\frac{U(f)}{\Phi_{0}^{2}(m_{\phi}^{2}-\omega_{0}^{2})^{2}}\right\}$ (8)
For concreteness we restrict most of our discussion to the sextic scalar potential studied in Ref. Heeck:2020bau . This can be parametrized as
$U(f)=\phi_{0}^{2}\left(\frac{m_{\phi}^{2}-\omega_{0}^{2}}{2}\,f^{2}(1-f^{2})^{2}+\frac{\omega_{0}^{2}}{2}\,f^{2}\right).$ (9)
III Spherical gauged Q-balls
We begin by reviewing the construction of spherical Q-balls. This section follows the analysis of Heeck:2021zvk .
The spherical Q-ball is approximately a sphere of radius $R$ with a constant internal density. More precisely, in the interior, the scalar field is approximately at the false vacuum solution $f=f_{+}$ (where $f_{+}$ is the nonzero solution of the equation 5) and at the Q-ball surface, the field rapidly transitions to the true vacuum $f=0$. The scalar profile is thus to a first approximation
$f^{(0)}=\left\{\begin{array}[]{c}f_{+}\qquad\rho<R\\ 0\qquad\rho>R\end{array}\right.$ (12)
The gauge field is a nondynamical field, and is determined from the scalar charge distribution through the Maxwell equation 6. For the scalar profile above, the solution for the gauge field is
$\alpha A^{(0)}=\left\{\begin{array}[]{cc}\Omega+Q_{in}{\sinh(\alpha f_{+}\rho)\over\rho}&~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\rho<R\\ {Q_{out}\over\rho}&~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\rho>R\end{array}\right.$ (15)
Note that the typical scale of variation of $A$ is inversely proportional to $\alpha$.
Near $\rho\sim R$, the scalar field is varying very quickly over a short distance of order 1, which we call the transition region. However, since the gauge field varies much more slowly than the scalar, we can take the gauge field to be approximately smooth over the transition region. Continuity of the gauge field and its derivative at $\rho=R$ then implies
$Q_{in}=-{\Omega\over{\alpha f_{+}\cosh(\alpha f_{+}R)}}\qquad\qquad Q_{out}=R\Omega\left(1-{1\over{\alpha f_{+}R}}{\tanh(\alpha f_{+}R)}\right)$ (16)
We note that the value of the gauge field at $\rho=R$ is
$\alpha A_{R}^{(0)}=\Omega\left(1-{1\over\alpha f_{+}R}{\tanh(\alpha f_{+}R)}\right)$ (17)
We can now require consistency between the obtained gauge field and the initial ansatz for the scalar profile. To do this, we solve the equation for the scalar field in the gauge field background (15). This equation is now
${d^{2}f^{(0)}\over d\rho^{2}}+\frac{2}{\rho}{df^{(0)}\over d\rho}=f^{(0)}\left(1-4(f^{(0)})^{2}+3(f^{(0)})^{4}+{\Omega_{0}^{2}}-\left(\Omega-\alpha A^{(0)}\right)^{2}\right)$ (18)
This is particularly important to solve in the transition region, where the scalar field varies quickly from the false vacuum to the true vacuum. This requires a very precise relation between the potential parameters and the radius of the Q-ball, as discussed in Heeck:2021zvk for the case of global Q-balls, and which we now review.
In the gauged Q-balls case that we are discussing here, the scalar field in the transition region would satisfy the equation
${d^{2}f^{(0)}\over d\rho^{2}}+\frac{2}{R}{df^{(0)}\over d\rho}=f^{(0)}\left(1-4(f^{(0)})^{2}+3(f^{(0)})^{4}+{\Omega_{0}^{2}}-(\Omega_{R}^{(0)})^{2}\right)$ (19)
where we have replaced the explicit factor of $\rho$ by $R$, and set the gauge field to be its value at $\rho=R$. This has allowed us to replace $\left(\Omega-\alpha A^{(0)}\right)$ by the constant
$\Omega_{R}^{(0)}=\left(\Omega-\alpha A^{(0)}_{R}\right)={\Omega~{}{\tanh(\alpha f_{+}R)}\over\alpha f_{+}R}$ (20)
This maps the problem to the form of a global Q-ball equation, and we can use the methods in  Heeck:2021zvk to solve it.
We can treat the term $\frac{2}{R}{df^{(0)}\over d\rho}$ as a small frictional term. In its absence, there is an exactly conserved energy, and this allows the equation to be solved exactly; consistency of energy conservation implies that this is a solution in the limiting case when the false vacuum and the true vacuum have the same energy. In this limit, the solution of the equation (19) is
$f^{(0)}(\rho)=\frac{f_{+}}{\sqrt{1+2e^{2(\rho-R)}}}\,.$ (21)
If the false vacuum and the true vacuum have slightly different energies, the frictional term must be included if the solution is to exist, and furthermore $R$ must have a precise value. Specifically, the change in the energy between the two vacua should be balanced by the effect of the frictional term. The details of the derivation of the resulting consistency condition are in Heeck:2021zvk and show that the radius must satisfy
$((\Omega_{R}^{(0)})^{2}-\Omega_{0}^{2})R=1+{\cal O}({1\over R})$ (22)
This can be shown Heeck:2021zvk to be an accurate prediction for the Q-ball radius.
IV Perturbed Q-balls
We now look for potential instabilities of these Q-balls.
We do these by considering deformations of the spherical Q-ball. The deformation of the Q-ball will in general not lead to a solution of the equations of motion. The deformations may grow, indicating that the Q-ball is unstable, or they may decay away. However, if there is a transition between stable Q-balls and unstable Q-balls, then exactly at threshold, the deformation will neither grow nor decay, and therefore the deformed Q-ball will be a solution of the equations of motion.
Now the spherical Q-ball has an approximately constant density in the interior, and zero density outside. We will consider perturbations of the Q-ball where the density is assumed to be approximately the same constant in the interior, but where the surface deforms in some way. It is likely that other deformations (e.g. where the internal density is nonconstant) are possible; we leave the search for these to future work. The assumptions we make will hence lead to a conservative estimate for the region of stability (i.e. our analysis will not find the most general instability, but any Q-balls that we find to be unstable will indeed be unstable).
It is fairly immediate that under these assumptions on the perturbations, a spherical perturbation would change the charge. We must therefore consider perturbations with an angular dependence, such that the total charge does not change to leading order.
At this point, then, we are led to consider a perturbation where the Q-ball surface is deformed to be at
$\rho=R(\theta)\equiv R(1+\epsilon P_{L}(\cos(\theta))).$ (23)
$P_{L}(\cos(\theta))$ is a Legendre polynomial; we shall denote this as $P_{L}$, suppressing the $\theta$ dependence. $\epsilon$ parametrizes the size of the perturbation, and can be taken to be parametrically small.
We now follow the same approach as the previous section to construct these Q-balls.
First, we will make an initial ansatz that the scalar field is at the false vacuum in the interior and zero outside i.e.
$f=\left\{\begin{array}[]{c}f_{+}\qquad\rho<R(1+\epsilon P_{L})\\ 0\qquad\rho>R(1+\epsilon P_{L})\end{array}\right.$ (26)
The equation of motion 6 for the gauge field can then be solved to find
$\alpha A=\left\{\begin{array}[]{c}\Omega+Q_{in}{\sinh(\alpha f_{+}\rho)\over\rho}+\epsilon P_{in}F_{L}(\alpha f_{+}\rho)P_{L}\qquad\rho<R(1+\epsilon P_{L})\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}Q_{out}{1\over\rho}+\epsilon P_{out}{1\over\rho^{L+1}}P_{L}\qquad~{}~{}~{}~{}~{}~{}~{}~{}\rho>R(1+\epsilon P_{L})\end{array}\right.$ (29)
where $F_{L}(\alpha f_{+}\rho)=j_{L}(i\alpha f_{+}\rho)$, and $j_{L}$ is a spherical Bessel function. We list the first few $F_{L}$ in an appendix.
We again require the gauge field and its derivative to be continuous at the boundary. This yields
$Q_{in}=-{\Omega\over{\alpha f_{+}\cosh(\alpha f_{+}R)}}\qquad\qquad Q_{out}=R\Omega\left(1-{1\over{\alpha f_{+}R}}{\tanh(\alpha f_{+}R)}\right)$ (30)
$P_{in}={\Omega\tanh(\alpha fR)\over\alpha fF^{\prime}_{L}(\alpha f\rho)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}\qquad\qquad P_{out}={\Omega R^{L+1}\tanh(\alpha fR)F_{L}(\alpha fR)\over F^{\prime}_{L}(\alpha fR)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}$ (31)
At the boundary, the value of the gauge field is now
$\alpha A_{R}=\Omega\left(1-{1\over\alpha f_{+}R}{\tanh(\alpha f_{+}R)}\right)+\epsilon P_{L}\Omega\left({F_{L}(\alpha fR)\tanh(\alpha fR)\over F^{\prime}_{L}(\alpha f\rho)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}-1+{\tanh(\alpha fR)\over\alpha fR}\right)$ (32)
We can now require again consistency between the obtained gauge field and the initial ansatz for the scalar profile. To do this, we solve the equation for the scalar field in the gauge field background (29). We focus on the transition region, where the scalar field varies quickly from the false vacuum to the true vacuum. Once again we note that the gauge field is slowly varying as a function of $\rho$, and can be taken to have a value $A_{R}$ over the transition region.
The differential equation for the scalar is
${\partial^{2}f\over\partial\rho^{2}}+\frac{2}{\rho}{\partial f\over\partial\rho}+{1\over\rho^{2}\sin\theta}\partial_{\theta}(\sin\theta\partial_{\theta}f)=f\left(1-4f^{2}+3f^{4}+{\Omega_{0}^{2}}-\Omega_{R}^{2}\right)$ (33)
where $\Omega_{R}=\Omega-\alpha A_{R}$. Since $f$ is now dependent on $\theta$, we have kept the angular derivatives in the Laplacian.
This is now a differential equation in two variables, which is obviously much more complex. However, we note that all the radial and angular dependence in the ansatz (26) is at the surface of the Q-ball, and hence these dependences are related. To make this precise, note that the ansatz (26) only depends on $\rho-R\epsilon P_{L}$. We therefore have
${1\over\rho^{2}\sin\theta}\partial_{\theta}(\sin\theta\partial_{\theta}f)={df\over d\rho}{1\over R}L(L+1)\epsilon P_{L}$ (34)
where we have only kept the terms linear in $\epsilon$. The angular derivative terms therefore yield terms proportional to ${df\over d\rho}$.
In the transition region, we can also set $\rho=R(\theta)=R(1+\epsilon P_{L}(\cos(\theta)))$. The scalar equation then becomes
${\partial^{2}f\over\partial\rho^{2}}+\left(\frac{2}{R(\theta)}+{1\over R}L(L+1)\epsilon P_{L}\right){\partial f\over\partial\rho}=f\left(1-4f^{2}+3f^{4}+{\Omega_{0}^{2}}-\Omega_{R}^{2}\right)$ (35)
The partial differential equation now only has derivatives with respect to $\rho$. This can then be solved separately for each $\theta$. For example, one can fix $\theta$, and then find the profile by starting at $f=f_{+}$ at $\rho=0$ and stepping outward to large $\rho$. We therefore get a profile for each value of $\theta$.
For each value of $\theta$, the equation is again of the form of a global Q-ball equation. We can write the equation (35) as
${d^{2}f\over d\rho^{2}}+\frac{2}{R_{eff}}{df\over d\rho}=f\left(1-4f^{2}+3f^{4}+{\Omega_{0}^{2}}-\Omega_{R}^{2}\right)$ (36)
where we have defined
$\frac{2}{R_{eff}}=\frac{2}{R(\theta)}+\epsilon{1\over R}L(L+1)P_{L}$ (37)
This is identical to the equation (19)in the previous section with the replacement $\Omega^{(0)}_{R}\to\Omega_{R},R\to R_{eff}$. The solution may therefore be read off from (22) and states that the perturbed Q-ball satisfies
$\Omega_{R}^{2}-{\Omega_{0}^{2}}={1\over R_{eff}}+{\cal O}({1\over R^{2}_{eff}})$ (38)
This may be simplified by matching terms linear in $\epsilon$ on both sides, and yields the equation for a stable deformed Q-ball as
${\Omega F_{L}(\alpha fR)\tanh(\alpha fR)\over F^{\prime}_{L}(\alpha f\rho)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}-\Omega\left(1-{\tanh(\alpha fR)\over\alpha fR}\right)=-{(L+2)(L-1)\over 4\Omega^{(0)}_{R}R}$ (39)
This is also the relation that gives the instability threshold. That is, if there is a transition from a region of stable Q-balls to a region of unstable Q-balls, the boundary is set by this relation.
What if this relation is not satisfied? Then the deformed Q-ball does not satisfy the equations of motion, and it will either relax back to the spherical Q-ball, or the deformation grows, indicating an instability. At least close to the threshold, we can track this by giving $\epsilon$ a small time dependence. Following the same derivation, we find
${R\ddot{\epsilon}\over\epsilon}=-4\Omega\Omega^{(0)}_{R}\left({F_{L}(\alpha fR)\tanh(\alpha fR)\over F^{\prime}_{L}(\alpha f\rho)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}-\left(1-{\tanh(\alpha fR)\over\alpha fR}\right)\right)-{(L+2)(L-1)\over R}$ (40)
Therefore, if the right-hand side of the above equation is negative, the Q-ball is stable, but if it is positive, there is a instability.
V Comparison with Numerics
We will now compare this theoretical prediction for the instability threshold of gauged Q-balls to the numerical results of  Kinach:2022jdx .
The authors of Kinach:2022jdx consider Q-ball instabilities in several types of potentials; one of the models they consider is a gauged Q-ball in a sextic potential with
$V=|\phi|^{2}-{1\over 2}|\phi|^{4}+{0.2\over 3}|\phi|^{6}$ (41)
For this potential they present three results:
(i) For a particular parameter point P2 (details below) the Q-ball is unstable. The instability has $L=4$.
(ii) For a particular parameter point P1 (details below) the Q-ball is stable.
(iii) For a particular choice of parameters, they find the instability threshold.
We review these results, and compare to our prediction.
(i) The parameter point P2 has $\omega=0.9958,e=0.02,r=65$. After rescaling to match our conventions, we find that this parameter point corresponds to
$\Omega_{0}^{2}=1/15\qquad\Omega=1.028\qquad R=62.9\qquad\alpha=0.057\qquad\Omega^{(0)}_{R}=0.2587$ (42)
For these parameters, we calculate the right-hand-side of eqn (40) for each value of $L$. We find that the right-hand-side vanishes for $L=1$; this reflects the translational mode of the Q-ball, and must be ignored. The right-hand-side is positive for $L=2,3,4$ and hence all these modes are predicted to lead to an instability. The value of the RHS for $L=5$ is negative, but very close to zero, and this result may be changed in a more accurate analysis. For $L>5$, the right-hand-side is negative, and hence these modes do not lead to an instability .
This is to be compared with the results of Kinach:2022jdx , which finds that the Q-ball with this parameter choice is unstable, with an instability mode with $L=4$. They also find the instability is near the surface of the Q-ball. This is all consistent with our results.
(ii) The parameter point P1 discussed in Kinach:2022jdx has $e=0.17,\omega=0.9976$. This corresponds in our notation to $\Omega=1.029,\alpha=0.481$. The value of $\Omega_{0}$ is again $1/\sqrt{15}$.
For such large values of $\alpha$, the analytical prediction of the previous section cannot be trusted. Nevertheless, we can see what the formula predicts. We find that the lowest instability threshold (i.e. the smallest unstable $R$) is predicted to occur for $L=2$ at a value $\alpha R=6.7$. This corresponds to a predicted threshold radius of about 14 (that is, Q-balls with a radius smaller than 14 are predicted to be stable, while larger Q-balls are predicted to be unstable).
The Q-ball radius at parameter point P1 is not given in Kinach:2022jdx , but we may compare the charges at these two parameter points (the charge for P1 is given as 387.5 while the charge for P2 is given as $1.480\times 10^{6}$) to estimate the ratio of the raidii for these two parameter points. This shows that the radius at parameter point P1 should be about 4. Our prediction would therefore be that the Q-ball at parameter point P1 is stable. This is in agreement with the numerical results.
(iii) For $e=0.02$ (corresponding to $\alpha=0.057$ in our conventions), the authors of Kinach:2022jdx also present the value of the instability threshold i.e. the first value of $R$ for which the Q-ball is unstable. This is found to be very close to the radius where ${d\Omega\over dR}$ becomes positive. This is at $\alpha R=0.51$.
Let us see what the equation (40) predicts. Unfortunately, we do not have the other parameter values of the numerical calculation, and we are forced to make estimations of these. Specifically, we need to find $\Omega,\Omega^{(0)}_{R}$ at the threshold.
We first calculate $\Omega^{(0)}_{R}$ using (22) to find
$\Omega^{(0)}_{R}=\left(\sqrt{1/15+{1\over R^{2}}}\right)$ (43)
For $\Omega$, we use the known values at parameter point P2, as well as equation (20) to find $\Omega$ as
${\Omega\over 1.028}={\Omega_{R}^{(0)}\over 0.2587}{\alpha f_{+}R_{t}\over{\tanh(\alpha f_{+}R_{t})}}{\tanh(3.26)\over 3.26}$ (44)
Substituting $\Omega,\Omega^{(0)}_{R}$ into equation (40) gives us an equation determining the instability threshold for each $L$. We find that the smallest radius of instability occurs for $L=2$, and for a predicted value of $\alpha R=1.99$. We therefore predict that all radii above this value are unstable.
Here we have a significant difference between our predicted instability threshold of $\alpha R=1.99$, and the numerically calculated threshold Kinach:2022jdx of $\alpha R=0.51$. We discuss this further in the next section.
VI Discussion and Conclusion
We have constructed an explicit perturbation of a gauged Q-ball, and shown that in some parameter ranges, the equations of motion directly show that these perturbations can be unstable. These perturbations have angular dependence, and are localized near the surface of the Q-ball. We have found an explicit prediction for the condition on the radius (and other parameters) for the modes to become unstable.
This is, to our knowledge, the first calculation that shows analytically that perturbations with angular dependence can lead to an instability in gauged Q-balls.
We have compared our predictions to the numerical calculation in Kinach:2022jdx , and shown that we correctly predict the instability/stability of the two parameter points analyzed in that paper. Furthermore, it was found in the numerical analysis that one of the points had an unstable mode with $L=4$, and our formula is able to reproduce this by showing that the $L=4$ mode is indeed unstable. Another parameter point was found to be stable, and this is reproduced by our formula. We also find that increasing the gauge coupling makes the Q-balls more stable, a surprising result which was also found in the numerics.
Finally, we note that our predicted boundary between stable and unstable parameters does not agree perfectly with the results of Kinach:2022jdx . It appears that any Q-ball found to be unstable by our formula is also found to be unstable in the numerical calculations; however, there appear to be Q-balls which are found numerically to be unstable which are not captured by our analysis. This suggests strongly that there are other perturbations that we have missed that will extend the region of instability. On a related note, the region of instability in the numerical results appears to be linked to the condition ${d\omega\over dQ}>0$, at least for small gauge couplings. We do not see such a link, which is again suggestive that a different form of perturbation may be relevant for Q-balls. This analysis is left for future work.
VII Acknowledgments
This work was supported in part by NSF Grant No. PHY-2210283. We thank M. Smolyakov for a comment on an initial version of this paper.
VIII Appendix A: Radial solutions
For completeness, we list the first few $F_{L}$ solutions used in the paper
$F_{0}$ $={\sinh x\over x}$ (45)
$F_{1}$ $=\sinh x\left({1\over x^{2}}\right)-\cosh x\left({1\over x}\right)$ (46)
$F_{2}$ $=-\cosh x\left({3\over x^{2}}\right)+\sinh x\left({1\over x}+{3\over x^{3}}\right)$ (47)
$F_{3}$ $=-\cosh x\left({1\over x}+{15\over x^{3}}\right)+\sinh x\left({6\over x^{2}}+{15\over x^{4}}\right)$ (48)
$F_{4}$ $=-\cosh x\left({10\over x^{2}}+{105\over x^{4}}\right)+\sinh x\left({1\over x}+{45\over x^{3}}+{105\over x^{5}}\right)$ (49)
$F_{5}$ $=-\cosh x\left({1\over x}+{105\over x^{3}}+{945\over x^{5}}\right)+\sinh x\left({15\over x^{2}}+{420\over x^{4}}+{945\over x^{6}}\right)$ (50)
$F_{6}$ $=-\cosh x\left({21\over x^{2}}+{1260\over x^{4}}+{10395\over x^{6}}\right)+\sinh x\left({1\over x}+{210\over x^{3}}+{4725\over x^{5}}+{10395\over x^{7}}\right)$ (51)